# Interpretable LLM-based Table Question Answering

### v2

@@ Results.
   - Atomic Operation 76.65% (-0.09) 71.74% (-0.19)
   - NL Plan 67.96% (-8.78) 66.20% (-5.73)
   - SQL 64.19% (-12.55) 66.54% (-5.39)
We conclude that all components (atomicity, planning, and SQL) contribute meaningfully to POS ’s improved interpretability, but SQL execution stands out as the key driver. By translating each atomic step into an explicit SQL query, POS naturally produces step-by-step attribution maps that pinpoint exactly which rows and columns influence the final answer. This clarity is especially valuable for verification: if a highlighted cell in the explanation is irrelevant or obviously mismatched to the question, a user or another LLM can readily infer that the model’s final prediction is suspect. The plan itself also matters significantly, presumably because it makes the chain-of-reasoning easier to follow. Enforcing atomic operations is beneficial, but their impact is more modest. This suggests that while simple transformations help reduce complexity—and improve model accuracy (see Table 6 for POS vs. Text-to-SQL accuracy)—the bulk of the interpretability gain comes from explicit programmatic execution (SQL) and a structured plan.
Table 6. Accuracy (%) for TabFact and WikiTQ using GPT-3.5 and GPT-4o-mini. “Breakdown” indicates whether queries are decomposed into sub-problems (see Figure 21). “Transformed by” refers to whether intermediate tables are transformed by an LLM or a program (see Figure 223). “Answered by” specifies whether the final answer is generated by an LLM or a program(see Figure 24). LLM-only approaches provide the final answer without table transformations. Bold values indicate the best performance for each model and dataset.
Method Accuracy (%) Breakdown Tables transformed by Final answer by
@@
4.2. Evaluating Table QA Performance
Baselines
We compare POS with several baseline methods, categorizing them into three groups based on how table transformation and answer generation are performed: LLM-only, program-only, and hybrid approaches. Unless otherwise noted, we use a temp = 00 and top-p = $1$ for sampling.
LLM-only.
Program-only.
Program-based approaches generate explicit programs to perform table transformation and answer the question. Latent Program Algorithm (LPA) (Chen et al., 2020) frames TabFact verification as a program synthesis task, converting input queries into sequential operations (e.g., min, max, count, filter) executed via Python-Pandas. On the other hand, Text-to-SQL (Rajkumar et al., 2022) translates a natural language query directly into a single SQL command, which is then applied to the input table to generate the answer.
Hybrid.
Hybrid approaches combine the strengths of LLM reasoning and programs to perform Table QA and achieve state-of-the-art accuracy. DATER (Ye et al., 2023) uses an LLM to extract relevant sub-tables, while breaking queries into sub-queries and executing SQL commands to retrieve factual information. Binder (Cheng et al., 2023) takes a different approach by converting natural language questions into executable programs. It blends API calls with symbolic language interpreters like SQL or Python to address reasoning gaps that cannot be handled through offline methods alone. Lastly, CoTable (Wang et al., 2024) dynamically plans a sequence of predefined table operations–such as selecting rows or adding columns, allowing it to iteratively transform the table based on the intermediate information. Despite their differences, they all share a common strategy: they input the final simplified table along with the original query into an LLM to produce the final answer.
4.2.1. Table QA for TabFact and WikiTQ
As shown in Table 6, POS delivers strong performance on both TabFact and WikiTQ across two different LLM backbones. When paired with GPT-3.5, POS achieves 78.31% accuracy on TabFact and 54.80% on WikiTQ, yielding substantial gains over LLM-only methods such as End-to-End QA, Few-Shot QA, and Chain-of-Thought. It also surpasses other program-only baselines by wide margins: for instance, POS outperforms Text-to-SQL by +13.6 points and LPA by +9.41 points on TabFact.
Using GPT-4o-mini, POS reaches 82.70% accuracy on TabFact and 59.32% on WikiTQ, surpassing DATER and obtaining the highest WikiTQ accuracy among tested approaches. Although Binder is slightly stronger on TabFact (84.63%), POS remains highly competitive against these hybrid methods. Furthermore, by generating explicit programs, POS offers unique interpretability advantages: users can verify each transformation step and gain deeper insights into the model’s reasoning pipeline, as discussed in Section 4.1.3.
These results demonstrate the great performance of POS across both GPT-3.5 and GPT-4o-mini, consistently outperforming LLM-only and most program-only methods while remaining competitive with hybrid approaches.
Table 7. Accuracy of POS across varying table sizes. The Pearson correlation reveals negligible relationships between table size and accuracy, highlighting the robustness of POS.
@@ 4.2.1. Table QA for TabFact and WikiTQ
4.2.2. Performance vs. table size analysis
A natural concern for real-world Table QA systems is whether performance degrades as tables grow larger and more complex. To address this, we provide a quantitative evaluation of POS on TabFact and WikiTQ, stratified by table size. Specifically, following the methodology in Wang et al. (2024), we sort each dataset by token count (our proxy for table size) and split it into three bins of equal size (small, medium, and large). We use GPT4o-mini as the backbone LLM and report accuracy in Table 7.
On TabFact, POS consistently maintains 79–85% accuracy across varying table sizes, with negligible correlation ($-0.006$) to length. WikiTQ shows a modest drop for larger tables ($-0.023$ correlation, 46.7–56.1%).
The robustness of POS is particularly meaningful given that previous methods—such as Binder, DATER, and CoTable—experience significant accuracy drops with large tables (i.e., tables longer than 4K tokens; see Table 3 in  Wang et al. (2024)). For instance, Binder and DATER can see accuracy declines of 30–50 percentage points, while CoTable exhibits a 25-percentage point decrease. This degradation is primarily due to the challenges LLMs face when processing long contexts. In contrast, POS’s SQL-based executions remain unaffected by table length, making it well-suited for real-world scenarios where table sizes may scale to millions of tokens.
4.2.3. SQL vs. LLM-based table transformations.
We investigate the impact of removing SQL execution from POS in Section 4.1.6. Without SQL, not only does interpretability decline, but accuracy on the TabFact and WikiTQ datasets also suffers. Specifically, we ask the LLM to transform the table directly rather than executing SQL. We find that this leads to a negligible decrease in Table QA accuracy on TabFact but a substantial drop on WikiTQ, indicating that relying on the black-box reasoning of LLMs for table transformations can severely impact model accuracy. We argue that this is likely due to the likelihood of LLM hallucinations or errors when handling complex tables (Chen, 2023; Wang et al., 2024). Additionally, this approach diminishes interpretability, as the transformations of the table are no longer transparent or traceable, which is quantified by Table 5.
4.3. Efficiency Analysis
Table 8. Efficiency analysis on WikiTQ. SC indicates self-consistency usage, LLM shows total (and breakdown) of LLM calls, and DB is the number of database queries. POS requires fewer LLM/database calls while achieving better efficiency than other methods.
@@ 4.3. Efficiency Analysis
Efficiency in Table QA is crucial because reducing the number of LLM calls and database queries directly lowers computational costs and improves scalability. We use the efficiency benchmark introduced in Chain-of-Table (Wang et al., 2024), which measures the number of LLM calls required to answer a WikiTQ question. Additionally, we propose using the number of database queries (i.e., table transformations) that reflects the computational workload on the table database.
POS employs highly deterministic atomic steps during planning, eliminating the need for the costly self-consistency prompting (Wang et al., 2022) typically required by Binder, DATER, and CoTable. As a result, POS requires only four LLM calls per question—significantly fewer than Binder (50), DATER (100), or CoTable (25). Regarding database queries, POS is also more efficient than others with only two queries per question, compared to five of CoTable, 50 of Binder, or 20 of DATER.
4.4. POS for Free-form Table QA
POS processes Table QA queries end-to-end through SQL commands, with table transformations executed offline on the input table and without access to external knowledge. This SQL-only pipeline restricts POS’s applicability to tasks requiring creativity, such as generating free-form answers. To address this, we extend POS to the free-form Table QA task (FeTaQA) (Nan et al., 2022) by integrating an LLM call in the final step to generate free-form natural language answers, following the algorithm illustrated in Figure 2.
We compare POS with End-to-End and Few-Shot QA and find that POS consistently outperforms both methods (see Table 9). This improved performance (+0.98 points in BLEU and +1.83 points in ROUGE-L compared to Few-Shot QA) is attributed to intermediate SQL executions of POS, which retrieve fine-grained and relevant information to generate answers. In contrast, End-to-End and Few-Shot QA process the entire input table at once, making it challenging for the model to pinpoint and make use of the correct information.
Table 9. Evaluation on FeTaQA with GPT-4o-mini.
@@ 4.4. POS for Free-form Table QA
Please note that we compare POS against these two end-to-end baselines—selected specifically to showcase the adaptability of POS in generating free-form natural language responses while preserving its interpretability advantages (in intermediate table transformations). As we prioritize interpretability, we have not optimized POS accuracy in free-form table QA, and therefore we do not include many hybrid QA baselines in  Table 9.
5. Discussion
In this section, we examine several natural questions regarding the robustness and interpretability of POS.
How do POS handle real-world “messy” tables? We acknowledge that truly “wild” tables—e.g., those with nested or repeated headers—are common in real-world scenarios. While we have adopted cleaned datasets from prior works (Wang et al., 2024; Ye et al., 2023) to facilitate SQL executions, we have not yet evaluated our approach on highly noisy table data. To address this, future works can integrate a wrapper that normalizes input tables (Nahid and Rafiei, 2024a) prior to running POS.
How to detect SQL code-generation errors, such as syntax issues or ambiguous references? Current version of POS incorporates a fallback mechanism: if an error occurs during query processing, a LLM is invoked to resolve the query using an end-to-end QA approach. This strategy follows with common practices in Table QA (Ye et al., 2023; Cheng et al., 2023; Wang et al., 2024). To avoid relying on fallback mechanisms, we propose implementing a proactive error detection and correction pipeline that validates each generated SQL query against the table schema. When an error is detected, the model automatically regenerates the query using the execution history log as context, thereby reducing error rates and minimizing the need for fallback.
What are the most common error types encountered by POS? We observed that errors in POS often originate from the planning stage rather than from the Step-to-SQL. For example, the planner often omits necessary condition checks in its atomic steps. As LLMs continue to improve their planning capabilities, we expect POS to benefit accordingly in terms of performance while retaining its strong interpretability. Additionally, SQL’s exact-matching behavior can lead to inherent issues, such as failing to match ‘bjørn’ with ‘bjon’. Please refer to Appendix I for illustrative examples.
@@ 5. Discussion
Is POS completely interpretable? No; no AI system is entirely interpretable. In computer vision, for example, methods such as concept bottleneck models provide human-understandable concepts (Ismail et al., [n.d.]) for model decisions, but the reasoning underlying the network to generate the concepts remains unintelligible. Likewise, large language models can generate chain-of-thought rationales, but the internal hidden states driving these thoughts cannot be directly understood by humans (Wei et al., 2022; Madsen et al., 2024).
POS bottlenecks the decision-making into human-interpretable concepts—namely, natural-language steps that are translated into SQL. These SQL statements deterministically determine the answer, allowing users to follow the logical chain of transformations on the table. Although the underlying LLM may involve complex internal processes (in generating natural-language steps), our focus is on making the model’s decision-making fully interpretable, following the approach of prior interpretability works (Taesiri et al., 2022; Ismail et al., [n.d.]).
6. Conclusion
We introduce Plan-of-SQLs or POS, an interpretable, effective, and efficient method for LLM-based Table QA. By decomposing queries into atomic steps, POS ensures that each transformation is simple, traceable, and easy to understand. This approach provides users with step-by-step reasoning for the generated answers, addressing a critical gap in existing LLM-based Table QA models, where explanations are often ambiguous. We expect that as LLMs continue to improve, POS will inherently become more accurate while maintaining its interpretability.
7. Limitations
First, using programmatic execution for table transformation, we share a common limitation with existing works (also baselines in Table 6(Ye et al., 2023; Cheng et al., 2023; Wang et al., 2024) where we use fallback when an intermediate operation fails. We observe that a small subset of samples cannot be processed end-to-end using SQL (3.16% for TabFact and 13.58% for WikiTQ). Since operations often fail due to special symbols in the input table, adopting an advanced table preprocessing approach (Nahid and Rafiei, 2024a) can effectively solve the problem. We also take an exploratory approach by planning one step at a time that improves accuracy (see Appendix I). Second, our work does not yet investigate the interpretability in free-form Table QA (FeTaQA) where we need to rely on the black-box reasoning of LLMs to generate the final answer. A potential solution to improve interpretability in this case is to create hybrid explanations that combine POS-generated explanations for intermediate steps with grounding (Hendricks et al., 2018) for the final step.
Appendix

### v3

@@ 4.1.6 Ablation study on POS interpretability
   — Atomic operations 76.65% (-0.09) 71.74% (-0.19)
   — NL planning 67.96% (-8.78) 66.20% (-5.73)
   — SQL execution 64.19% (-12.55) 66.54% (-5.39)
By translating each atomic step into an explicit SQL query, POS naturally produces step-by-step attribution maps that pinpoint exactly which table cells influence the final answer. This clarity is especially valuable for verification: if a highlighted entry in the explanation is irrelevant or obviously mismatched to the question, a user can readily infer that the model’s final answer is suspect. The natural-language plan itself also matters significantly, presumably because it makes the chain-of-reasoning much easier to follow. Although ablating atomicity causes only a minor drop—likely because modern LLMs already employ strong step-by-step reasoning priors (Wei et al., 2022)—we argue that explicit atomic steps remain meaningful for human understanding. Please refer to Appendix D for qualitative examples.
Table 6: Accuracy (%) for TabFact and WikiTQ using GPT-3.5 and GPT-4o-mini. “Decomposed” indicates whether queries are decomposed into sub-problems (Fig. 21). “Transformed by” refers to whether intermediate tables are transformed by an LLM or a program (Fig. 223). “Answered by” specifies whether the final answer is generated by an LLM or a program (Fig. 24). LLM-only approaches provide the final answer without table transformations. The best performance for each model and dataset is shown in bold.
Method Accuracy (%) Decomposed Tables transformed by Final answer by
@@
4.2 Evaluating Table QA Performance
Baselines.
We compare POS with several baseline methods, categorizing them into three groups based on how table transformation and answer generation are performed: LLM-only, program-only, and hybrid approaches. Unless otherwise noted, we use a temp = 00 and top-p = $1$ for LLM generation.
4.2.1 Table QA for TabFact and WikiTQ
As shown in Tab. 6, POS achieves strong performance on both TabFact and WikiTQ on two different LLM backbones. When paired with GPT-3.5, POS achieves 78.31% accuracy on TabFact and 54.80% on WikiTQ, yielding substantial gains over LLM-only methods such as End-to-End QA, Few-Shot QA, and Chain-of-Thought. It also surpasses other program-only baselines by wide margins: for instance, POS outperforms Text-to-SQL by +13.6 points and LPA by +9.41 points on TabFact. Compared to hybrid approaches, which combine LLM and program-based operations, POS offers a compelling alternative. Although POS scores lower accuracy on both benchmarks than these hybrid methods, its exclusive use of program-based transformations ensures complete transparency of each reasoning step, making it much easier for users to verify and understand the underlying decision-making process.
Using GPT-4o-mini, POS reaches 82.70% on TabFact and 59.32% on WikiTQ. While hybrid approaches like Binder and CoTable score higher on TabFact, POS achieves the best performance on WikiTQ. This demonstrates that with a more advanced language model, POS remains competitive while offering great interpretability, as shown in Sec. 4.1.3.
Table 7: Accuracy of POS across varying table sizes. The Pearson correlation reveals negligible relationships between table size and accuracy.
@@ 4.2.1 Table QA for TabFact and WikiTQ
4.2.2 Performance vs. table size analysis
A natural concern for Table QA systems is whether the performance degrades as input tables grow larger and more complex. To answer this, we provide a quantitative evaluation of POS accuracy on TabFact and WikiTQ, stratified by table size. Specifically, following the methodology in Wang et al. (2024), we sort each table by token count (our measure for table size) and split datasets into three bins of equal size (small, medium, and large). We use GPT-4o-mini as the backbone LLM and report the QA accuracy in Tab. 7.
We observe that POS maintains stable accuracy on TabFact (79.1 $\rightarrow$ 85.2 $\rightarrow$ 81.5%, r=–0.006), though on WikiTQ performance drops from 56.1% (small) to 46.7% (medium) before recovering slightly to 47.8% (large) ($<$10 percentage points; r=–0.023). This suggests POS is resilient on some benchmarks but can suffer on tables of intermediate size in more complex tasks.
We find that POS is more robust than existing methods—such as Binder, DATER, and CoTable—which all suffer significant accuracy drops with large tables (i.e., tables longer than 4K tokens; see Table 3 in Wang et al. (2024)). For instance, Binder and DATER experience accuracy declines of 30–50 percentage points, dropping to 6.41% and 34.62%, respectively, while CoTable degrades to 44.87% on WikiTQ. This degradation is primarily due to the challenges LLMs face when processing long contexts. In contrast, POS’s SQL-based executions remain resilient to variations in table length, making it well-suited for real-world scenarios where table sizes may scale to millions of tokens.
4.3 Efficiency Analysis
Table 8: Efficiency analysis on WikiTQ. SC denotes self-consistency usage, LLM represents the total number of LLM calls (detailed in Breakdown), and DB is the number of database queries. Notably, POS requires up to 25$\times$ fewer calls/queries than other methods.
@@ 4.3 Efficiency Analysis
Efficiency in Table QA is crucial because reducing the number of LLM calls and database queries directly lowers computational costs and improves scalability. We leverage the efficiency benchmark introduced by Wang et al. (2024), which measures the number of LLM calls required to answer a WikiTQ question. Additionally, we propose using the number of database queries (i.e., table transformations) that reflects the computational workload on the table database.
As POS employs highly deterministic and atomic steps, it eliminates the need for the costly self-consistency prompting (Wang et al., 2022) required by Binder, DATER, and CoTable. As a result, POS requires only four LLM calls per question—significantly fewer than CoTable (25), Binder (50), and DATER (100). Regarding database queries, POS is also more efficient than others with only two queries per question, compared to five of CoTable, 50 of Binder, or 20 of DATER.
4.4 POS for Free-form Table QA
POS processes Table QA queries end-to-end using SQL commands, performing table transformations exclusively on the input table and without accessing to external knowledge. This SQL-only pipeline restricts POS’s applicability to tasks requiring creativity, such as generating paragraph-like answers. To address this, we extend POS to the free-form Table QA task (FeTaQA) (Nan et al., 2022) by integrating an LLM call in the final step to generate free-form natural language answers, following the algorithm illustrated in Fig. 2.
Table 9: Results on FeTaQA, a free-form Table QA task (using GPT-4o-mini as the base LLM). POS outperforms both 0-shot end-to-end and few-shot QA.
@@ 4.4 POS for Free-form Table QA
We compare POS with End-to-End and Few-Shot QA and find that POS consistently outperforms both methods (see Tab. 9). This improved performance (+0.98 points in BLEU and +1.83 points in ROUGE-L compared to Few-Shot QA) is attributed to intermediate SQL executions of POS, which retrieve fine-grained and relevant information to generate final answers. In contrast, End-to-End and Few-Shot QA process the entire input table at once, making it challenging for the model to pinpoint and make use of the correct data.
@@
Please note that we compare POS against these two LLM-only baselines—selected specifically to showcase the adaptability of POS in generating free-form natural language responses while preserving its interpretability advantages (in intermediate table transformations). As we prioritize interpretability, we have not optimized POS accuracy in free-form Table QA, and therefore we do not include many hybrid QA baselines in Tab. 9.
5 Discussion and Future Works
We address pertinent questions regarding the robustness and interpretability of POS; more are in Appendix C.
How does POS handle real-world “messy” tables? We have so far evaluated on benchmark tables that are semi-structured and relatively clean (as preprocessed by Wang et al. (2024) and Ye et al. (2023)). In practice, tables “in the wild” may have merged cells, nested headers, irregular formats, or other noise. POS in its current form does not explicitly tackle such noise. A practical extension would be to integrate a preprocessing step that normalizes and restructures messy tables (e.g., using a tool like NormTab (Nahid & Rafiei, 2024a)) before applying POS. This could allow POS to maintain high accuracy and interpretability even on more complex, non-canonical tables.
How can POS detect errors in its own SQL generation? POS uses a fallback that could diminish the program-only nature: if SQL execution fails, an LLM answers the question end-to-end—triggered in 3.16% of TabFact and 13.58% of WikiTQ samples, similar to other Table QA methods (Ye et al., 2023; Cheng et al., 2023; Wang et al., 2024). Tab. 6 reports POS accuracy using this fallback mechanism. To minimize such fallbacks, we propose implementing a proactive error detection and correction mechanism that validates each generated SQL query against the table schema. When an error is detected, the model regenerates the query using the execution history log as context.
@@ 5 Discussion and Future Works
Is POS completely interpretable? No; no AI system is entirely interpretable. In computer vision, for example, methods such as concept bottleneck models provide human-understandable concepts (Ismail et al., 2023) for model decisions, but the reasoning underlying the network to generate the concepts remains unintelligible. Likewise, LLMs can generate chain-of-thought rationales, but the internal hidden states driving these thoughts cannot be directly understood (Wei et al., 2022; Madsen et al., 2024). POS bottlenecks the decision-making into human-interpretable operations—namely, natural-language steps that are translated into SQLs. They deterministically produce the answer, allowing users to follow the logical chain of table transformations. Although the underlying LLM may involve black-box internal processes (in generating natural-language steps), our focus is on making model’s decision-making interpretable, aligning with prior and current interpretability research (Taesiri et al., 2022; Ismail et al., 2023; Zhao et al., 2024).
Can every table-based question be decomposed into a set of atomic steps? Yes; by design POS breaks any question over a table into a sequence of atomic operations—filter, select, group, aggregate, sort, limit, join, set operations, and so on—each implemented by a single, standalone SQL statement. Because SQL is relationally complete (i.e., it can express every relational-algebra query), this set of primitives (or atomic operations) can express any query, from simple counts to complex multi-table analytics [^2] https://www.cwblogs.com/posts/relational-algebra/.
The only limitation arises for truly free-form queries that demand narrative or summary generation (e.g., “Summarize the basic information of the football clubs in Saint Petersburg.”). In such cases, POS still applies atomic SQL steps to extract and prepare the data, then defers only the final writing step to a language model, preserving full transparency of all intermediate transformations.
What are the most common failure modes of POS? We observed that errors in POS mostly originate from the planning stage rather than from the Step-to-SQL. For example, the planner often omits necessary condition checks in its atomic steps or performs steps in the orders that lead to the wrong answers (refer to Sec. H.1.3 for the importance of step order). As LLMs continue to improve their planning capabilities (Huang et al., 2022), we expect POS to benefit accordingly in terms of performance while retaining its strong interpretability. Please refer to Appendix J for examples and analysis of POS errors.
Why POS explanations do not contain SQLs? We designed POS to present natural-language steps rather than raw SQL because, in our user studies, including SQL in the explanations tended to overwhelm users. The natural language descriptions are equivalent to the SQL commands—thanks to the atomicity constraint, which guarantees minimal-to-no discrepancy between the two—yet they are far more readable and accessible. This design choice strikes a balance between interpretability and clarity, ensuring that users grasp the model’s reasoning without being burdened by technical SQL syntax.
6 Conclusion
We introduce Plan-of-SQLs (POS), an interpretable, effective, and efficient approach to Table QA with large language models. POS decomposes a table question into simple atomic steps and executes each step using SQL, thereby ensuring that each transformation in the reasoning process is transparent. This design fills a critical gap in current LLM-based Table QA models, which often produce answers without clear explanations. Our experiments demonstrate that prioritizing interpretability does not come at the expense of performance: POS achieves explanation quality superior to existing methods (as confirmed by both human evaluations and LLM judges) and competitive accuracy on Table QA benchmarks. Moreover, POS accomplishes this with an order-of-magnitude fewer LLM calls and database queries than prior approaches, making it a more efficient and scalable solution for real-world use. We also find that human evaluators and LLM judges largely agree on their evaluations of explanations, indicating that LLM-based automated evaluation is feasible. The only significant limitation we observe is rooted in the LLM’s planning capability—if the LLM produces a suboptimal plan, POS can err, which aligns with recent observations by Zhao et al. (2024). Encouragingly, as LLMs become more powerful in planning, we expect the performance of POS to naturally improve while preserving its inherent interpretability. In summary, POS represents a step toward Table QA models that not only deliver accurate answers but also provide human-understandable explanations.
Acknowledgement
GN was supported by an Auburn University Presidential Graduate Research Fellowship. AN was supported by NSF Grant #2145767, NAIRR award #240116, donations from NaphCare Foundation, and research gifts from Adobe Research.
Disclaimer
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates ("JP Morgan”) and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
Appendix for:
Interpretable LLM-based Table Question Answering